# Boosting Prediction with Data Missing Not at Random

### v1

@@ 6.4 Simulation Results
Figure 1: Prediction results in simulation studies in the MAR scenario. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the values for S-MAE and S-RMSE, respectively.
Figure 2: Prediction results in simulation studies in the MNAR scenario. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the values for S-MAE and S-RMSE, respectively.
To visualize predicted values derived from different methods, Figure 3 shows boxplots of the average predicted values, ${n_{2}}^{-1}\sum^{n_{2}}_{i=1}\hat{f}^{*}_{n_{1}}(X_{i})$, across 500 simulations, along with the expected value $E(Y_{i})$, whose expression is provided in the Supplementary Materials. Under MAR, all methods yield similar results, though the N method tends to deviate more from the R method than the proposed methods. However, with MNAR in Setting 1, the R and proposed methods (i.e., IPW, IPWN, and BJ) produce fairly close results, while the N methods exhibit noticeable bias. In Setting 2, with a large number of correlated covariates, IPWN does not perform satisfactorily, as the MNAR estimation procedure of Morikawa and Kim (2021) assumes independent covariates.
Figure 3: Prediction results in simulation studies: Boxplots of ${n_{2}}^{-1}\sum^{n_{2}}_{i=1}\hat{f}^{*}_{n_{1}}(X_{i})$ obtained from the five methods in combination with the three loss functions, for 500 simulations, where the grey line indicates the value of $E(Y_{i})$. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the MAR and MNAR settings, respectively.
Finally, we assess the convergence of the proposed boosting methods. The results confirm the convergence of our boosting algorithm, as long as estimates of $\pi_{2}(y,x)$, defined in the Supplementary Materials, using the method of Morikawa and Kim (2021) are available; further details are provided in the Supplementary Materials. We note that the method proposed by Morikawa and Kim (2021) is specifically designed for MNAR data and may not perform well for MAR data, particularly in scenarios with high-dimensional covariates. In such cases, divergence can occur frequently, as acknowledged by Morikawa and Kim (2021), thus preventing the estimation of $\pi_{2}(y,x)$ and halting the boosting procedure. This issue may stem from various factors related to starting values for the nonlinear optimization function, the complexity of the optimization process, and the large number of parameters.
@@ 6.5 Model Misspecification
Figure 6: Performance under model misspecification: Boxplots of $n^{-1}\sum^{n}_{i=1}\hat{f}^{*}_{n}(X_{i})$ obtained from the eight methods in combination with the three loss functions, for 500 simulations, where the grey line indicates the value of $E(Y_{i})$. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the MAR and MNAR settings, respectively.
@@ 7 Analysis of KLIPS Data
To illustrate the proposed methods, we apply them to analyze a dataset from the Korean Labor and Income Panel Survey (KLIPS), which contains demographics information for 2501 Korean workers. The KLIPS data are frequently analyzed in the MNAR literature, including Kim and Yu (2011), Wang et al. (2014), Shao and Wang (2016), Morikawa et al. (2017), and Morikawa and Kim (2021).
@@
where $\gamma_{j}$ is the parameter for $j=0,1,\ldots,4$. Given the missing indicator $R_{i}=1$ and covariates $X_{i,1}$, $X_{i,2}=j$, $X_{i3}=k$ for $j=1,2,3$ and $k=1,2$, the response $Y_{i}$ follows $N(\beta_{0,j,k}+\beta_{1,j,k}X_{i,1}+\beta_{2,j,k}X_{i,1}^{2}+\beta_{3,j,k}X_{i,1}^{3}+\beta_{4,j,k}X_{i,1}^{4},\sigma_{j,k}^{2})$, with parameters $\beta_{j,k}=(\beta_{0,j,k},\beta_{1,j,k},\beta_{2,j,k},\beta_{3,j,k},\beta_{4,j,k})^{\top}$ and $\sigma_{j,k}$. Following Morikawa and Kim (2021), for each combination of $j$ and $k$, we use the stepwise AIC to select the best model.
We analyze the data using the methods considered in Section 6.3 except the R method. Let $\hat{f}^{*}_{n}(X_{i})$ denote the predicted value for the observed covariates $X_{i}$ with $i=1,\ldots,n$, obtained from each method. To visualize those estimates, for each estimate $\hat{f}^{*}_{n}(\cdot)$, we fix the second argument to be 1, 2, or 3, and the third argument to be 1 or 2, and then in Figure 7, we plot $\hat{f}^{*}_{n}(\cdot)$ against the first argument, denoted $X_{1}$. The results yielded from the proposed methods are similar and tend to deviate from those produced by the N method at the left and right ends of the first argument’s range. Additionally, in Figure 8, we plot boxplots for the predicted values $\left\{\hat{f}^{*}_{n}(X_{i}):i=1,\ldots,n\right\}$ obtained from all considered methods.
Figure 7: Plots of the predicted values $\hat{f}^{*}_{n}(X_{i})$ versus its first argument $X_{i1}$ for $i=1,\ldots,n$, obtained from the proposed methods and the N method, with the second argument of $\hat{f}^{*}_{n}(X_{i})$ fixed as 1, 2, or 3 (indicated by the left to right columns, respectively) and with the third argument of $\hat{f}^{*}_{n}(X_{i})$ fixed as 1 or 2 (indicated by the top to bottom rows, respectively).
Figure 8: Boxplots of $\left\{\hat{f}^{*}_{n}(X_{i}):i=1,\ldots,n\right\}$ obtained from the proposed methods and the N method.
@@ 8 Discussion
Supplementary Materials
Text document:
A .pdf file contains regularity conditions, proofs of Propositions 1 and 2, Theorems 1 and 2, and additional simulation results.
R-code:
A .zip file, named “code.zip”, includes R scripts for the simulations and data analyses presented in Sections 6 and 7.
Acknowledgments
The authors thank the review team for their comments on the initial version of the manuscript and Dr. Morikawa for providing the KLIPS data and the R code used in Morikawa and Kim (2021). Yi is a Canada Research Chair in Data Science (Tier 1). Her research is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Canada Research Chairs Program.

### v2

@@ 6.4 Simulation Results
Figure 1: Prediction assessments in the MAR scenario. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the values for S-MAE and S-RMSE, respectively.
Figure 2: Prediction assessments in the MNAR scenario. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the values for S-MAE and S-RMSE, respectively.
To visualize predicted values derived from different methods, Figure 3 shows boxplots of the average predicted values, ${n_{2}}^{-1}\sum^{n_{2}}_{i=1}\hat{f}^{*}_{n_{1}}(X_{i})$, across 500 simulations, along with the expected value $E(Y_{i})$, whose expression is provided in the Supplementary Materials. Under MAR, all methods yield similar results, though the N method tends to deviate more from the R method than the proposed methods. However, with MNAR in Setting 1, the R and proposed methods (i.e., IPW, IPWN, and BJ) produce fairly close results, while the N methods exhibit noticeable bias. In Setting 2, with a large number of correlated covariates, IPWN does not perform satisfactorily, as the MNAR estimation procedure of Morikawa and Kim, (2021) assumes independent covariates.
Figure 3: Boxplots of ${n_{2}}^{-1}\sum^{n_{2}}_{i=1}\hat{f}^{*}_{n_{1}}(X_{i})$ obtained from the five methods in combination with the three loss functions, where the dashed line indicates the value of $E(Y_{i})$. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the MAR and MNAR settings, respectively.
Finally, we assess the convergence of the proposed boosting methods. The results confirm the convergence of our boosting algorithm, as long as estimates of $\pi_{2}(y,x)$, defined in the Supplementary Materials, using the method of Morikawa and Kim, (2021) are available; further details are provided in the Supplementary Materials. We note that the method proposed by Morikawa and Kim, (2021) is specifically designed for MNAR data and may not perform well for MAR data, particularly in scenarios with high-dimensional covariates. In such cases, divergence can occur frequently, as acknowledged by Morikawa and Kim, (2021), thus preventing the estimation of $\pi_{2}(y,x)$ and halting the boosting procedure. This issue may stem from various factors related to starting values for the nonlinear optimization function, the complexity of the optimization process, and the large number of parameters.
@@ 6.5 Model Misspecification
Figure 6: Performance under model misspecification: Boxplots of $n^{-1}\sum^{n}_{i=1}\hat{f}^{*}_{n}(X_{i})$ obtained from the eight methods in combination with the three loss functions, where the dashed line indicates the value of $E(Y_{i})$. Top and bottom rows correspond to Settings 1 and 2, respectively; left and right columns correspond to the MAR and MNAR settings, respectively.
@@ 7 Analysis of KLIPS Data
To illustrate the proposed methods, we apply them to analyze a dataset from the Korean Labor and Income Panel Survey (KLIPS), which contains demographics information for 2501 Korean workers. The KLIPS data are frequently analyzed in the MNAR literature, including Kim and Yu, (2011), Wang et al., (2014), Shao and Wang, (2016), Morikawa et al., (2017), Morikawa and Kim, (2021), and Bian et al., 2025b .
@@
where $\gamma_{j}$ is the parameter for $j=0,1,\ldots,4$. Given the missing indicator $R_{i}=1$ and covariates $X_{i,1}$, $X_{i,2}=j$, $X_{i3}=k$ for $j=1,2,3$ and $k=1,2$, the response $Y_{i}$ follows $N(\beta_{0,j,k}+\beta_{1,j,k}X_{i,1}+\beta_{2,j,k}X_{i,1}^{2}+\beta_{3,j,k}X_{i,1}^{3}+\beta_{4,j,k}X_{i,1}^{4},\sigma_{j,k}^{2})$, with parameters $\beta_{j,k}=(\beta_{0,j,k},\beta_{1,j,k},\beta_{2,j,k},\beta_{3,j,k},\beta_{4,j,k})^{\top}$ and $\sigma_{j,k}$. Following Morikawa and Kim, (2021), for each combination of $j$ and $k$, we use the stepwise AIC to select the best model.
We analyze the data using the methods considered in Section 6.3 except the R method. Let $\hat{f}^{*}_{n}(X_{i})$ denote the predicted value for the observed covariates $X_{i}$ with $i=1,\ldots,n$, obtained from each method. To visualize those estimates, for each estimate $\hat{f}^{*}_{n}(\cdot)$, we fix the second argument to be 1, 2, or 3, and the third argument to be 1 or 2, and then in Figure 7, we plot $\hat{f}^{*}_{n}(\cdot)$ against the first argument, denoted $X_{1}$. The results yielded from the proposed methods are similar and tend to deviate from those produced by the N method at the left and right ends of the first argument’s range. Additionally, in Figure 8, we plot boxplots for the predicted values $\left\{\hat{f}^{*}_{n}(X_{i}):i=1,\ldots,n\right\}$ for all considered methods.
Figure 7: The predicted values $\hat{f}^{*}_{n}(X_{i})$ versus its first argument $X_{i1}$ for $i=1,\ldots,n$, obtained from the proposed methods and the N method, with the second argument of $\hat{f}^{*}_{n}(X_{i})$ fixed as 1, 2, or 3 (left to right columns) and with the third argument of $\hat{f}^{*}_{n}(X_{i})$ fixed as 1 or 2 (top to bottom rows).
Figure 8: Boxplots of $\left\{\hat{f}^{*}_{n}(X_{i}):i=1,\ldots,n\right\}$.